OpenStudy (anonymous):

amy paid $79.23 for a pair of running shoes during a 35% off sale what was the regular price? My answer is $121.89 reg. price is this correct

OpenStudy (anonymous):

$(79.23*100)/(100-35)=$121.89 reg. yes u are right

OpenStudy (anonymous):

solve then graph 5x>=15 I know the answer is 3 is the correct solution x<3 or x>=3

OpenStudy (anonymous):

since she got a discount of 35%, it means that 79.23 is 65% of the original price so let the regular price = x 65% of x = 79.23 65 ----- * x = 79.23 100 79.23 * 100 x = ----------- = 121.92 65 so the regular price of the shoes was $ 121.89

OpenStudy (anonymous):

harkirat can you help with this i am checking my ans. soybean meal is 16% protein, cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320-lb mixture that is 10% protein? so how many pounds of the cornmeal should be in the mixture? how many pounds of the soybean meal shoud be in the mixture?

OpenStudy (anonymous):

for previous question x>=3 becoz 5x = 5*3 = 15 so if x >=3 then 5x>=15

OpenStudy (anonymous):

now i will solve the soyabean one....

OpenStudy (anonymous):

let the amount of soyabean be = x lb then amount of cornmeal = (320-x)lb 1 lb of soyabean is 16% protein i.e 16/100 = 0.16 lb protein so x lbs of soyabean has = 0.16 * x = 0.16x lb of protein 1 lb of cormeal is 8% protein i.e. 8/100 = 0.08 lb protein so (320 - x) lbs of cornmeal has = 0.08 * (320-x) = ( 25.6 - 0.08x) lb of protein 1 lb of mixture is 10% protein i.e. 10/100 = 0.1 lb protein so 320 lbs of cornmeal has = 0.1 * 320 lb of protein ATQ, 0.16x + (25.6 - 0.08x) = 0.1 * 320 => 0.08x + 25.6 = 32 => 0.08x = 32 - 25.6 => x = 6.4/0.08 = 80 so there shud b 80 lbs of soyabean in the mixture and 320-80 = 240 lbs of cornmeal in the mixture

OpenStudy (anonymous):

my GOOD ANSWER clicks pls !!1 ( :))

OpenStudy (anonymous):

so 80 is for cornmeal and 32 for soybean

OpenStudy (anonymous):

No, 80 lbs of soyabean and 240 lbs of cornmeal !!!!